OpenStudy (anonymous):

Simplify: -(-1+5)^2 - 2(-4)

OpenStudy (anonymous):

-8. I just typed exactly that into my calculator :)

OpenStudy (geometry_hater):

First distribute the negative and the -2

OpenStudy (anonymous):

Hm, I got 18

OpenStudy (geometry_hater):

hold on a sec ill check it out

OpenStudy (anonymous):

ok

OpenStudy (markrijckenberg):

Follow the Order of Operations rules and you should find the right answer. There was a question about something similar on Facebook recently and more than 70% of the online population got it wrong....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (markrijckenberg):

So what is the answer? :-)

OpenStudy (anonymous):

33

OpenStudy (anonymous):

is that correct?

OpenStudy (markrijckenberg):

No, I believe the answer is -8. -(-1+5)^2 - 2(-4) = -4^2+8= -16+8=-8